himig:

In a flower garden, there are 5 tulips for every 8 daisies. If there are 45 tulips, how many daisies are there?

XxAngelGirlxX:

54 daisies?

jimthompson5910:

Let x be the number of daisies that go with 45 tulips. x is some positive whole number \(\Large \frac{\text{5 tulips}}{\text{8 daisies}}=\frac{\text{45 tulips}}{\text{x daisies}}\) \(\Large \frac{5}{8}=\frac{45}{x}\) \(\large 5x=8*45 ... \text{ cross multiply}\) I'll let you finish up